OpenStudy (laronjaxon):

simplfy √1 ________________ (4-3i)-(2+5i)

OpenStudy (mertsj):

First combine the like terms in the denominator

OpenStudy (laronjaxon):

(4-2)-(3I+5I)?

OpenStudy (anonymous):

2-8i

OpenStudy (mertsj):

Now multiply numerator and denominator by the conjugate of 2-8i which is 2+8i

OpenStudy (mertsj):

\[\frac{1}{2-8i}\times \frac{2+8i}{2+8i}\]

OpenStudy (laronjaxon):

-4+i ______ 4????/

OpenStudy (anonymous):

\[2+8i /4+64=2+8i/68=1/34+2/17i\]

OpenStudy (mertsj):

\[\frac{2+8i}{4+64}=\frac{2+8i}{68}=\frac{1}{34}+\frac{2}{17}i\]

OpenStudy (mertsj):

Yep. Good job.

OpenStudy (anonymous):

;)